OpenStudy (anonymous):

(x-1)y'+2xy=0 , y(2) = a : Solution?

OpenStudy (anonymous):

change it into : y'/y = -2x / (x-1) now integrate : -2x/(x-1) = -2 - 2/(x-1) so : ln|y| = -2x -2ln|x-1| + c y = C * e^(-2x -2ln|x-1|) can you do it now ?